in the formula c=sqrt of(2a squared+b squared). a=4 and b=5.my answer was 7.550? is this correct.
Answer by mukhopadhyay(490)

c=sqrt of(2a squared+b squared)
=>c=sqrt(2*4^2 + 5^2)
=>c=sqrt(2*16 + 25)
=>c=sqrt(32+25)
=>c=sqrt(57)
Exact answer of c is sqrt(57)
Approximate answer is what you have got.